### v2

@@ 7.3. Monads
    $L\subset(\Psi\circ\Phi)[K]\iff\exists M\subset\Phi[K]\;\bigl{(}L\subset\Psi[M]\bigr{)},$
    hence $T(\Psi\circ\Phi)=T(\Psi)\circ T(\Phi)$; Thus $T$ is indeed a functor.
(2) Naturality of $\eta_{X}$. For each $x\in X$, $\eta_{X}(x)=\{x\}$ is a continuous map (singletons are closed in compact Hausdorff spaces and form a continuous embedding into $\mathcal{K}(X)$), and hence upper hemicontinuous and compact-valued.
(3) Naturality of $\nu_{X}$. Given $\mathcal{A}\in\mathcal{K}(\mathcal{K}(X))$, $\nu_{X}(\mathcal{A})=\bigcup\mathcal{A}$ is continuous in the Vietoris topology, [^14] For basic opens $\langle U_{1},\dots,U_{m}\rangle\subset\mathcal{K}(X)$ we have $\nu_{X}^{-1}\!\langle U_{1},\dots,U_{m}\rangle=\bigl{\langle}\,\langle U_{1}\rangle,\dots,\langle U_{m}\rangle,\,\langle U_{1}\cup\cdots\cup U_{m}\rangle\bigr{\rangle},$ which is open in $\mathcal{K}(\mathcal{K}(X))$, see Definition 29. and hence upper hemicontinuous and compact-valued.
(4) Monad laws.
  • Left unit: $\nu_{X}\circ T(\eta_{X})(K)=\bigcup_{x\in K}\eta_{X}(x)=\bigcup_{x\in K}\{x\}=K$.
  • Right unit: $\nu_{X}\circ\eta_{T(X)}(K)=\nu_{X}(\{K\})=K$.
  • Associativity: $\nu_{X}\circ T(\nu_{X})=\nu_{X}\circ\nu_{T(X)}$ holds due to the associativity of set union.
Therefore, $(T,\eta_{X},\nu_{X})$ is a monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. ∎
@@
The monad in Theorem 21 is the classical Vietoris monad (Garner,, 2020), adapted to the category of compact Hausdorff spaces with compact-valued, upper hemicontinuous correspondences. The Eilenberg-Moore category $\mathbf{AlgV}$ of algebras for the Vietoris monad on compact Hausdorff spaces is equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous lattice homomorphisms (Day,, 1975; Wyler,, 1981).
@@
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$, $\mathsf{CRED}_{\alpha}$, and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}_{\alpha}=\kappa_{\alpha}$. That is, the Full CP Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$.
@@
We show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are compact-valued. The claim, then, follows from an argument similar to that in the Proof of Theorem 5. Throughout, $Y$ is compact Hausdorff by (i); Every closed subset of $Y$ is therefore compact.
1. $\kappa$. Fix $(\alpha,y^{n},\psi)$ with $\alpha\notin S_{n+1}$, so to satisfy (iii). Define $\beta(\alpha)\coloneqq\min\{g\in S_{n+1}:g>\alpha\}$ as in (8). The conformal transducer $\pi(y,y^{n},\psi)$ is an upper semicontinuous map of $y$: Every summand in its definition is the indicator of a closed set and finite averages preserve u.s.c. Hence
$\kappa(\alpha,y^{n},\psi)=\bigl{\{}y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\bigr{\}},$
a super-level set of a u.s.c. function, therefore closed in $Y$, and thus compact.
2. $\mathsf{CRED}$. For a countable generator of closed sets $\{A_{j}\}\subset\Sigma_{Y}$,
@@
$\mathsf{CRED}(y^{n},\psi)=\bigcap_{j}\bigl{\{}P\in\Delta_{Y}:P(A_{j})\leq\overline{\Pi}_{y^{n},\psi}(A_{j})\bigr{\}}.$
Since $P\mapsto P(A_{j})$ is weak-continuous and $j\mapsto\overline{\Pi}_{y^{n},\psi}(A_{j})$ is u.s.c. by Lemma 4, each half-space is closed in the compact space $\Delta_{Y}$; Intersection keeps closedness, and in turn compactness.
3. $\mathsf{IHDR}$. For fixed $(\alpha,\mathcal{M})$, put
@@
$C_{\alpha}(\mathcal{M})\coloneqq\bigl{\{}A\in\Sigma_{Y}:\underline{P}_{\!\mathcal{M}}(A)\geq 1-\alpha\bigr{\}}.$
Each $A\in C_{\alpha}(\mathcal{M})$ is closed (complements of open sets have lower probability $\leq\alpha$). Hence $\mathsf{IHDR}(\alpha,\mathcal{M})=\bigcap_{A\in C_{\alpha}(\mathcal{M})}A$ is an intersection of closed subsets of the compact $Y$, hence compact. ∎
@@
Theorem 22 gives us minimal sufficient conditions so that (i) Full Conformal Prediction is a morphism of a category, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, that has a monad, and (ii) the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. We now look at a subcategory of $\mathbf{WMeas}_{\text{uc}}$ that has a monad.
Definition 23 (Structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$).
Define a structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ as follows,
  • Objects: Compact Polish spaces $X$ equipped with their Borel $\sigma$-algebra $\mathcal{B}(X)$.
  • Morphisms: Uniformly compact-valued, weakly measurable correspondences $\Phi:X\rightrightarrows Y$.
It is immediate to see that structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is a subcategory of $\mathbf{WMeas}_{\text{uc}}$, as its objects need to be compact as well. We now show that it admits a monad.
Let $X$ be an object of $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Define an endofunctor $\mathcal{K}\colon\mathbf{WMeas}_{\text{uc}}^{\text{c}}\to\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, a unit transformation $\eta_{X}\equiv\eta:\mathrm{Id}_{\mathbf{WMeas}_{\text{uc}}^{\text{c}}}\!\Rightarrow\!\mathcal{K}$ and a multiplication $\nu_{X}\equiv\nu:\mathcal{K}\!\circ\!\mathcal{K}\Rightarrow\mathcal{K}$ by
$\mathcal{K}(X)$$\coloneqq\bigl{\{}C\subseteq X:C\neq\emptyset,\;C\text{ compact}\bigr{\}},$
$\mathcal{K}(\Phi)(C)$$\coloneqq\bigl{\{}\,\overline{\Phi[C]}\bigr{\}},\quad C\in\mathcal{K}(X),$
$\eta_{X}(x)$$\coloneqq\bigl{\{}\,\{x\}\bigr{\}},\quad x\in X,$
$\nu_{X}(\mathcal{C})$$\coloneqq\bigl{\{}\bigcup_{C\in\mathcal{C}}C\bigr{\}},\quad\mathcal{C}\in\mathcal{K}\!\bigl{(}\mathcal{K}(X)\bigr{)}.$
@@
Theorem 24 (Vietoris monad on $\mathbf{WMeas}_{\mathrm{uc}}^{\mathrm{c}}$).
The triple $(\mathcal{K},\eta,\nu)$ satisfies the monad axioms inside $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Moreover, $\mathcal{K}(X)\not\cong X$ for any $X$ other than a singleton, so the monad is non-trivial.
@@
    (Objects). For any compact Polish space $X$, $\mathcal{K}(X)$ endowed with the Vietoris topology is again compact Polish; Its Borel $\sigma$-algebra makes $\mathcal{K}(X)$ an object of the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$.
  2. (2)
    (Morphisms). Given $\Phi:X\rightrightarrows Y$ in $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, each fiber of $\mathcal{K}(\Phi)$ is a singleton $\{\overline{\Phi[C]}\}$, hence compact; The common compact bound is $K_{\Phi}$, the one already bounding $\Phi$. Weak measurability follows because inverse images of Vietoris subbasic open sets can be expressed via the measurable “hit” and “miss” sets of $\Phi$ (studied below). Functoriality is immediate,
    $\mathcal{K}(\operatorname{id}_{X})=\operatorname{id}_{\mathcal{K}(X)},\qquad\mathcal{K}(\Psi\circ\Phi)=\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi),$
    since closures commute with unions in compact spaces.
2. The unit $\eta$. For every $X$, $\eta_{X}$ has compact fibres $\{x\}$, is uniformly bounded by $\mathcal{K}(X)$, and is weakly measurable because $\eta_{X}^{u}(\langle U\rangle)\coloneqq\{x\in X:\eta_{X}(x)\cap\langle U\rangle\neq\emptyset\}=U\in\mathcal{B}(X)$ for each open $U\subseteq X$. Naturality of $\eta$ is trivial: $\mathcal{K}(\Phi)\circ\eta_{X}=\eta_{Y}\circ\Phi$ for every $\Phi:X\rightrightarrows Y$.
3. The multiplication $\nu$. For $\mathcal{C}\in\mathcal{K}\bigl{(}\mathcal{K}(X)\bigr{)}$, the union $\bigcup_{C\in\mathcal{C}}C$ is compact (union of a compact family of compact subsets in a compact space). Thus $\nu_{X}(\mathcal{C})\in\mathcal{K}(X)$, giving well-defined singleton fibres. Uniform boundedness is by $\mathcal{K}(X)$ itself. Weak measurability is checked on subbasic Vietoris open sets:
@@
  • Hit part: $\nu_{X}^{-1}\!\langle U\rangle=\bigl{\{}\mathcal{C}:\exists C\in\mathcal{C},\;C\cap U\neq\emptyset\bigr{\}},$ which is open in the Vietoris topology of $\mathcal{K}(\mathcal{K}(X))$.
  • Miss part: $\nu_{X}^{-1}\langle U_{1},\dots,U_{m}\rangle=\bigl{\{}\mathcal{C}:\textstyle\bigcup_{C\in\mathcal{C}}C\subseteq U_{1}\cup\dots\cup U_{m},\;C\cap U_{i}\neq\emptyset\text{ }\forall i\bigr{\}},$ again Vietoris-open.
Naturality of $\nu$ follows from $\nu_{Y}\circ\mathcal{K}(\mathcal{K}(\Phi))=\mathcal{K}(\Phi)\circ\nu_{X}$, since direct images commute with unions and closure.
4. Monad laws. For every $X$ the diagrams
$\mathcal{K}^{3}(X)\;\xrightarrow{\mathcal{K}\nu_{X}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)\quad=\quad\mathcal{K}^{3}(X)\;\xrightarrow{\nu_{\mathcal{K}(X)}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)$
and
$\mathcal{K}(X)\;\xrightarrow{\mathcal{K}\eta_{X}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)\quad$ $=\quad\mathcal{K}(X)\;\xrightarrow{\mathrm{id}}\;\mathcal{K}(X)$
$=\quad\mathcal{K}(X)\;\xrightarrow{\eta_{\mathcal{K}(X)}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)$
commute because unions of singletons reproduce the original compact set, and unions are associative.
5. Non-triviality. If $X$ has at least two points, then $\mathcal{K}(X)$ is strictly larger than $X$ under any continuous, measure-compatible identification; Hence $(\mathcal{K},\eta,\nu)$ is not isomorphic to the identity monad. ∎
The monad that we found on the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is again the classic Vietoris monad, adapted to the category of compact Polish spaces (with their Borel $\sigma$-algebra), with uniformly compact-valued, weakly measurable correspondences. It is immediate to see that Theorem 8 still holds for $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, since we simply require the extra assumption that the objects are compact measurable Polish spaces. In addition, the same consideration regarding category $\mathbf{AlgV}$ of algebras for the Vietoris monad on compact Hausdorff spaces that we presented after Theorem 21, holds here as well, since Polish spaces are always Hausdorff.
@@
Proposition 25 (Forgetful Embedding between $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ and $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$).
@@
$F\;:\;\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}\;\rightarrow\;\mathbf{WMeas}_{\text{uc}}^{\text{c}},$
@@
    For every metrizable compact Hausdorff space $X$, pick any compatible complete separable metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
@@
    $F(\Phi)\;\coloneqq\;\Phi\qquad(\text{same graph and fibres}).$
@@
Proof.
1. $F(\Phi)$ is a morphism in $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Because $Y$ itself is compact, $\Phi(x)\subseteq Y$ for all $x$, so $K_{\Phi}\coloneqq Y$ is a common compact bound; Thus $F(\Phi)$ is uniformly compact-valued. To show weak measurability, fix an open $O\subseteq Y$ and put $H\coloneqq\{x\in X:\Phi(x)\cap O\neq\emptyset\}.$ If $x_{0}\notin H$ then $\Phi(x_{0})\subseteq Y\setminus O$ (a closed set). Upper hemicontinuity yields a neighborhood $U$ of $x_{0}$ with $\Phi(U)\subseteq Y\setminus O$, so $U\subseteq X\setminus H$. Hence $X\setminus H$ is open and $H$ is closed, thus Borel; This is precisely the weak measurability condition for $F(\Phi)$.

### v3

@@ 6.3. Monads
$(\nu_{X}\circ T\eta_{X})(K)=\bigl{\{}\textstyle\bigcup_{C\in\mathcal{C}^{\prime}}C:\emptyset\neq\mathcal{C}^{\prime}\subset\eta_{X}[K],\mathcal{C}^{\prime}\in\mathcal{K}(\mathcal{K}(X))\bigr{\}}=\{C\in\mathcal{K}(X):C\subset K\},$
so $\mathrm{id}_{\mathcal{K}(X)}\subseteq\nu_{X}\circ T\eta_{X}$ (take $\mathcal{C}^{\prime}=\eta_{X}[K]$ to recover $K$), but equality need not hold, hence laxity.
Right unit (strict): $\eta_{\mathcal{K}(X)}(K)=\{K\}$ and therefore $(\nu_{X}\circ\eta_{\mathcal{K}(X)})(K)=K$, so $\nu\circ\eta T=\mathrm{id}$ holds as equality.
Associativity (lax): For $\mathscr{B}\in\mathcal{K}(\mathcal{K}(\mathcal{K}(X)))$, elements of $(T\nu_{X})(\mathscr{B})$ are compact families $\mathcal{A}\subset\{\bigcup_{A\in\mathcal{C}}A:\mathcal{C}\in\mathscr{B}\}$, and
$(\nu_{X}\circ T\nu_{X})(\mathscr{B})=\Bigl{\{}\bigcup_{A\in\mathcal{A}^{\prime}}A:\mathcal{A}^{\prime}\subset(T\nu_{X})(\mathscr{B}),\mathcal{A}^{\prime}\in\mathcal{K}(\mathcal{K}(X))\Bigr{\}}.$
Each such union is a union of unions of sets from members of $\mathscr{B}$; Regrouping gives a union of sets from $\bigcup_{\mathcal{C}\in\mathscr{B}}\mathcal{C}$, so
$\nu_{X}\circ T\nu_{X}\subseteq\nu_{X}\circ\nu_{\mathcal{K}(X)},$
establishing the lax associativity.
4) Restriction to CHaus. If we restrict morphisms to single-valued continuous maps $f:X\to Y$, then $T(f)(K)=f[K]$, $\eta$ and $\nu$ are the classical unit and multiplication, and the three lax inclusions above become equalities, recovering the usual Vietoris monad. ∎
The construction in Theorem 19 adapts the Vietoris hyperspace mechanism to the correspondence category $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$. Because morphisms are upper hemicontinuous, compact-valued correspondences, the hyperspace assignment $T$ is a strict endofunctor: For morphisms $\Phi,\Psi$ we have
$T(\Psi\circ\Phi)=T(\Psi)\circ T(\Phi).$
The key point is that for compact $K\subset X$, $\Phi[K]$ is compact, so $T(\Psi\circ\Phi)(K)=\{L:L\subset\Psi[\Phi[K]]\}=(T\Psi\circ T\Phi)(K)$. The resulting structure is lax only at the level of monad laws, where
$\mathrm{id}\subseteq\nu\circ T\eta,\qquad\nu\circ\eta T=\mathrm{id},\qquad\nu\circ T\nu\subseteq\nu\circ\nu T.$
@@
By contrast, on the classical category $\mathbf{CHaus}$ (compact Hausdorff spaces with continuous maps), the Vietoris hyperspace $X\mapsto\mathcal{K}(X)$ together with $T(f)(K)=f[K]$, $\eta(x)=\{x\}$, and $\mu(\mathcal{A})=\bigcup_{A\in\mathcal{A}}A$ defines a strict monad (Garner,, 2020). In that classical setting, the Eilenberg-Moore (EM) category of algebras for the Vietoris monad is equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous lattice homomorphisms Day, (1975); Wyler, (1981).
Hence, the EM-$\mathbf{ContLat}$ equivalence applies to the classical Vietoris monad on $\mathbf{CHaus}$. For our correspondence-based (lax) variant on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$, we would need a theory of lax algebras; We do not pursue this here.
@@
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$.
@@
Fix a nonempty subset $\mathscr{F}_{0}\subset\mathscr{F}$ that is closed, uniformly bounded, and equicontinuous on the compact domain $Y^{n}\times Y$. By Arzelà-Ascoli, $\mathscr{F}_{0}$ is compact in the sup-norm, hence $Y^{n}\times\mathscr{F}_{0}$ is compact Hausdorff. In what follows we restrict $\kappa_{\alpha}$ and $\mathsf{CRED}$ to $Y^{n}\times\mathscr{F}_{0}$.
We show that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}$ are compact-valued and have closed graphs; By Lemma 3 they are upper hemicontinuous and hence morphisms in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. Throughout, $Y$ is compact Hausdorff by (i); Thus, closed subsets of $Y$ are compact.
(1) $\kappa_{\alpha}$ is compact-valued. Fix $(\alpha,y^{n},\psi)$ with $\alpha\notin S_{n+1}$ and set $\beta(\alpha)=\min\{g\in S_{n+1}:g>\alpha\}$ (as in (8)). For each $i$, the set
$\{y\in Y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n},y)\}$
is closed in $y$ (difference of continuous functions $\geq 0$). Hence $y\mapsto\pi(y,y^{n},\psi)$ is upper semicontinuous as a finite average of indicators of closed sets, and
$\kappa_{\alpha}(y^{n},\psi)=\{y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\}$
is a superlevel set of a u.s.c. map, thus closed in $Y$, hence compact.
(2) $\mathsf{CRED}$ is compact-valued. We can write that
@@
$\mathsf{CRED}(y^{n},\psi)=\bigcap_{A\in\Sigma_{Y}\text{, }A\text{ closed}}\Bigl{\{}P\in\Delta_{Y}:P(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)\Bigr{\}}.$
For each closed $A$, the map $P\mapsto P(A)$ is weak-continuous on the compact Hausdorff space $\Delta_{Y}$, and $(y^{n},\psi)\mapsto\overline{\Pi}_{(y^{n},\psi)}(A)$ is upper semicontinuous by Lemma 4. Thus each half-space is closed in $\Delta_{Y}$, and arbitrary intersections of closed sets are closed; Hence $\mathsf{CRED}(y^{n},\psi)$ is closed in $\Delta_{Y}$, therefore compact.
(3) $\mathsf{IHDR}_{\alpha}$ is compact-valued. In this compact Hausdorff setting we adopt the closed-sets presentation
@@
$\mathsf{IHDR}_{\alpha}(\mathcal{M})\coloneqq\bigcap\bigl{\{}F\subseteq Y:F\text{ closed and }\underline{P}_{\mathcal{M}}(F)\geq 1-\alpha\bigr{\}}.$
Monotonicity of lower probabilities ensures this presentation agrees with the original one whenever $\mathcal{M}=\mathsf{CRED}(y^{n},\psi)$ and $\pi$ is consonant (by Caprio et al., 2025a (, Proposition 5), $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))=\{y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\}$, which is closed). By construction the right-hand side is an intersection of closed subsets of $Y$, hence closed; Since $Y$ is compact Hausdorff, it is compact.
Having established compact-valuedness, the same closed-graph arguments used in the proof of Theorem 5 (together with Lemma 3) give upper hemicontinuity of $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}$; Thus they are morphisms of $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. Finally, Caprio et al., 2025a (, Proposition 5) yields the equality $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))=\kappa_{\alpha}(y^{n},\psi)$ for all admissible $(\alpha,y^{n},\psi)$, i.e. the diagram commutes. ∎
@@
Theorem 20 gives us minimal sufficient conditions so that (i) Full Conformal Prediction is a morphism of a category, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, that has a monad, and (ii) the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. We now show that a subcategory of $\mathbf{WMeas}_{\text{uc}}$ admits a monad.
@@
Theorem 21 (Vietoris monad on $\mathbf{WMeas}_{\mathrm{uc}}$ (with hemicontinuity)).
For this theorem only, strengthen the morphism condition in $\mathbf{WMeas}_{\mathrm{uc}}$ by additionally requiring both upper and lower hemicontinuity (still uniformly compact-valued and weakly measurable). [^15] It is easy to see that this class is still closed under composition. Define the endofunctor $\mathcal{K}$, unit $\eta$, and multiplication $\nu$ by
$\mathcal{K}(X)$ $\coloneqq\{C\subseteq X:C\neq\emptyset,C\text{ compact}\}\text{with the Vietoris topology (Borel $\sigma$-algebra),}$
$\mathcal{K}(\Phi)(C)$ $\coloneqq\{L\in\mathcal{K}(Y):L\subset\overline{\Phi[C]}\},\qquad C\in\mathcal{K}(X),$
$\eta_{X}(x)$ $\coloneqq\bigl{\{}\{x\}\bigr{\}},$
$\nu_{X}(\mathcal{C})$ $\coloneqq\bigl{\{}\textstyle\bigcup_{C\in\mathcal{C}}C\bigr{\}},\qquad\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)).$
Then, $(\mathcal{K},\eta,\nu)$ is a lax monad on $\mathbf{WMeas}_{\mathrm{uc}}$ (under this added hemicontinuity standing assumption). Moreover, if $|X|>1$, the unit $\eta_{X}:X\to\mathcal{K}(X)$ is not an isomorphism (it is not surjective), so the monad is nontrivial.
@@
(Objects). If $X$ is compact Polish, then $\mathcal{K}(X)$ with the Vietoris topology is compact Polish; Its Borel $\sigma$-algebra makes $\mathcal{K}(X)$ an object of $\mathbf{WMeas}_{\mathrm{uc}}$.
(Morphisms). Let $\Phi:X\rightrightarrows Y$ be a morphism (uniformly compact-valued with common compact bound $K_{\Phi}\subset Y$, weakly measurable, and both upper and lower hemicontinuous).
Each fiber of $\mathcal{K}(\Phi)$ is the singleton $\{\overline{\Phi[C]}\}$, which lies in $\mathcal{K}(K_{\Phi})$. Hence, $\mathcal{K}(\Phi)$ is uniformly compact valued with uniform bound $\mathcal{K}(K_{\Phi})\subset\mathcal{K}(Y)$. By a standard hyperspace continuity result (Aubin and Frankowska,, 1990; Beer,, 1993), if $\Phi$ has compact values and is both upper and lower hemicontinuous, then the map
$H:\mathcal{K}(X)\longrightarrow\mathcal{K}(Y),\qquad H(C)\coloneqq\overline{\Phi[C]},$
is continuous for the Vietoris topology. Hence $\mathcal{K}(\Phi)$ is continuous (and therefore weakly measurable) with singleton compact fibers uniformly bounded by $K_{\Phi}$.
Functoriality. $\mathcal{K}(\mathrm{id}_{X})=\mathrm{id}_{\mathcal{K}(X)}$. For $\Psi:Y\rightrightarrows Z$ also hemicontinuous with compact values,
$\bigl{(}\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi)\bigr{)}(C)=\bigl{\{}\overline{\Psi[\overline{\Phi[C]}]}\bigr{\}},\qquad\mathcal{K}(\Psi\circ\Phi)(C)=\bigl{\{}\overline{(\Psi\circ\Phi)[C]}\bigr{\}}.$
We claim $\overline{\Psi[\overline{\Phi[C]}]}=\overline{(\Psi\circ\Phi)[C]}$. One inclusion is monotonicity: $\Psi[\Phi[C]]\subset\Psi[\overline{\Phi[C]}]$, hence $\overline{(\Psi\circ\Phi)[C]}\subset\overline{\Psi[\overline{\Phi[C]}]}$. For the reverse, let $z\in\overline{\Psi[\overline{\Phi[C]}]}$. Then, there are sequences $y_{m}\in\overline{\Phi[C]}$ and $z_{m}\in\Psi(y_{m})$ with $z_{m}\to z$. By compactness, pass to subsequences with $y_{m}\to y\in\overline{\Phi[C]}$. Since $\Psi$ is upper hemicontinuous with compact values, its graph is closed, so $(y,z)\in\mathrm{Gr}(\Psi)$, i.e. $z\in\Psi(y)$. Because $y\in\overline{\Phi[C]}$, there are $u_{k}\in\Phi[C]$ with $u_{k}\to y$. By lower hemicontinuity of $\Psi$ at $y$, for every neighbourhood $W$ of $z$ there is a neighbourhood $N$ of $y$ with $\Psi(x)\cap W\neq\emptyset$ for all $x\in N$; Hence for large $k$ we can pick $v_{k}\in\Psi(u_{k})\cap W$. Thus $W$ meets $\Psi[\Phi[C]]$ for every neighbourhood $W$ of $z$, proving $z\in\overline{\Psi[\Phi[C]]}=\overline{(\Psi\circ\Phi)[C]}$. Therefore $\mathcal{K}(\Psi\circ\Phi)=\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi)$.
2. The unit $\eta$ and multiplication $\nu$. Both $\eta_{X}$ and $\nu_{X}$ have singleton fibers, hence uniformly compact-valued. For weak measurability (indeed continuity), compute on subbasic opens
@@
$\eta_{X}^{-1}\langle U\rangle=U,\qquad\nu_{X}^{-1}\langle U\rangle=\bigl{\{}\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)):\exists C\in\mathcal{C},C\cap U\neq\emptyset\bigr{\}},$
and similarly for miss-type sets $[U]$; these are Vietoris-open. Naturality of $\eta$ and $\nu$ is immediate from $\mathcal{K}(\Phi)[\{x\}]=\{\overline{\Phi[x]}\}$ and $\mathcal{K}(\Phi)\bigl{[}\{\overline{\cup\mathcal{C}}\}\bigr{]}=\{\overline{\cup\mathcal{K}(\Phi)[\mathcal{C}]}\}$.
3. Monad laws. The unit and associativity hold in the lax sense:
$\mathrm{id}_{\mathcal{K}(X)}\subseteq\nu_{X}\circ\mathcal{K}(\eta_{X}),\qquad\nu_{X}\circ\eta_{\mathcal{K}(X)}=\mathrm{id}_{\mathcal{K}(X)},\qquad\nu_{X}\circ\mathcal{K}(\nu_{X})\subseteq\nu_{X}\circ\nu_{\mathcal{K}(X)}.$
Indeed, $\mathcal{K}(\eta_{X})(C)$ consists of all compact families of singletons drawn from $C$, so $\nu_{X}\circ\mathcal{K}(\eta_{X})(C)$ contains $C$ (taking the full family of singletons). For the right unit, $\eta_{\mathcal{K}(X)}(C)=\{\{C\}\}$ and $\nu_{X}$ returns $\{C\}$. For associativity, unions of (compact families of) unions yield no more than the union of all underlying members, giving the inclusion.
4. Non-triviality. If $|X|>1$, then $\eta_{X}:X\to\mathcal{K}(X)$ is not surjective (non-singleton compact sets are not in its image), hence no natural isomorphism $\mathcal{K}\cong\mathrm{Id}$ exists.
Thus $(\mathcal{K},\eta,\nu)$ is a lax monad on $\mathbf{WMeas}_{\mathrm{uc}}$ under the added hemicontinuity assumption. ∎
The construction in Theorem 21 is a version of classical Vietoris monad, adapted to our measurable setting: Objects are compact Polish spaces (with their Borel $\sigma$-algebras), and morphisms are uniformly compact-valued, weakly measurable correspondences. for the hyperspace construction and functoriality of $\mathcal{K}$, we strengthened morphisms to be both upper and lower hemicontinuous (a standard hypothesis ensuring continuity of the induced map on hyperspaces). This added assumption is only used to establish the (lax) monad; All results proved earlier for $\mathbf{WMeas}_{\mathrm{uc}}$—including Theorem 24—do not rely on the monad and remain valid verbatim.
Regarding Eilenberg-Moore algebras: On the category $\mathbf{CHaus}$ of compact Hausdorff spaces, the Vietoris monad (Garner,, 2020) has Eilenberg-Moore category equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous homomorphisms (Day,, 1975; Wyler,, 1981). Since compact Polish spaces are precisely the compact metrizable Hausdorff spaces, restricting this equivalence to metrizable carriers yields the full subcategory of $\mathbf{ContLat}$ consisting of those algebras whose Lawson topology is metrizable.
@@
Proposition 22 (Forgetful Embedding between $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ and $\mathbf{WMeas}_{\text{uc}}$).
@@
$F:\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}\rightarrow\mathbf{WMeas}_{\text{uc}},$
@@
    For every metrizable compact Hausdorff space $X$, pick any compatible (complete separable) metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
@@
    $F(\Phi)\coloneqq\Phi\qquad(\text{same graph and fibers}).$
@@
The functor $F$ in Proposition 22 is defined on the full subcategory of $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$ with metrizable objects, since $(X,\mathcal{B}(X))$ must be compact Polish to lie in $\mathbf{WMeas}_{\mathrm{uc}}$. When working within that metrizable subcategory (as we do here), the statement holds verbatim. The choice of compatible metric on $X$ does not affect $\mathcal{B}(X)$.
Proof of Proposition 22.
1. $F(\Phi)$ is a morphism in $\mathbf{WMeas}_{\text{uc}}$. Because $Y$ is compact, $\Phi(x)\subseteq Y$ for all $x$, so $K_{\Phi}\coloneqq Y$ is a common compact bound; Thus $F(\Phi)$ is uniformly compact-valued. To show weak measurability, fix an open $O\subseteq Y$ and define $\varphi_{O}(y)\coloneqq d_{Y}\bigl{(}y,(Y\setminus O)\bigr{)}$, where $d_{Y}$ is any compatible (complete) metric on $Y$. Set
$h_{O}(x)\coloneqq\sup_{y\in\Phi(x)}\varphi_{O}(y).$
Since $\Phi$ is upper hemicontinuous with compact values, Berge’s maximum theorem implies $h_{O}$ is upper semicontinuous; Hence
$\{x:\Phi(x)\cap O\neq\emptyset\}=\{x:h_{O}(x)>0\}$
is open (in particular, Borel). This is precisely the weak measurability condition for $F(\Phi)$.